OpenStudy (anonymous):

can you check if i got these answers right? 1. Find the sum of the series 5+0+(-5)+...+ (-130) A. -3500 B. -1750 C. 625 D. -1820 (B is my answer) 2. Find the sixth term of the geometric sequence 1/2, 1/3, 2/9,... A. 32/2187 B. 32/729 C. 8/81 D. 16/243 (i choose D) 3. Give the first four terms of the geometric sequence for which a_1=-6 and r=4 A. -2,2,6,10 B. -6,-24,-96, -384 C. -6, -3/2, -3/8, -3/32 D. -24,-96,-384,-1536 (i choose A) 4. Which of the following is a geometric mean of 9 and 81? A. 729 B. 45 C. -27 D. 36 (i choose A)

OpenStudy (anonymous):

2. 1/2,1/3,2/9,.... a=1/2 r=(1/3)/(1/2)=1/3*2/1=2/3 tn=ar^(n-1) n=6 calculate t6

OpenStudy (anonymous):

The correct answer will be D (16/243)

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

is everything correct then?

OpenStudy (anonymous):

first four terms are a1,a1*r,a1*r*r,a1*r*r*r,a1*r*r*r*r

OpenStudy (anonymous):

4. r/9=81/r r^2=9*81=729 r=27

OpenStudy (anonymous):

so -27 would be the best option? (C)

OpenStudy (anonymous):

or 729?

OpenStudy (anonymous):

-27

OpenStudy (anonymous):

r^2=729 r=+-27 Here r=-27

OpenStudy (anonymous):

oh ok. are the rest of the problems good then?

OpenStudy (anonymous):

3. -6,-6*4,-6*4*4,-6*4*4*4,-6*4*4*4*4 =-6,-24,-96,-384

OpenStudy (anonymous):

1. not clear with statement.

OpenStudy (anonymous):

oh ok. thank you for helping me and for being patient :)